# Stability of heat kernel bounds under pointed Gromov-Hausdorff convergence

### v1

@@ Abstract
We present a method for constructing a heat kernel on the pointed Gromov-Hausdorff limit space and demonstrate the stability of heat kernel estimates under this convergence. Furthermore, we establish the Mosco convergence of the associated energy forms along a subsequence.
@@ 1 Introduction
    Since the assumptions and conclusions in Theorem 1.2 are stable under a bi-Lipschitz change of metric, we may instead assume that each $d_{j}$ is (uniformly) bi-Lipschitz equivalent to a length metric, which can be ensured by quasi-convexity [HKST15, Proposition 8.3.12] or by the chain condition [KM20, Proposition A.1].
@@
We briefly outline the proof of Theorem 1.2. The Radon measure $m_{\infty}$ on limit space is derived from the tightness of the push-forward measures. To obtain the limit Dirichlet form $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$, we first expand the heat kernel of the part of $(\mathcal{E}_{n},\mathcal{F}_{n})$ on the balls $B_{n}(p_{n},R)$ with Dirichlet boundary conditions, using their eigenvalues and eigenfunctions (see (4.1)). We then apply a version of Arzelà–Ascoli theorem (Lemma 5.1) to take the limit and obtain the limits of eigenvalues and eigenfunctions, which gives a heat kernel on $B_{\infty}(p_{\infty},R)$. Letting $R\to\infty$, we obtain a heat kernel on $X_{\infty}$ (Theorem 6.5). This heat kernel defines a semigroup (Theorem 6.6), corresponding to a Dirichlet form that satisfies the desired properties. The verification of Mosco convergence relies on the strong convergence of semigroups, as stated in the Mosco-Kuwae-Shioya’s theorem (Theorem 7.3).
@@
In [KS03, Section 5], the authors show that, under the uniform Poincaré inequality and other geometric conditions, the Dirichlet forms converge in the Mosco topology. However, it remains unknown whether the Poincaré inequality $\mathrm{(PI)}$ is stable—i.e., whether it holds for the limit Dirichlet form. In [HKST15, Section 11.6], the stability of the $p$-Poincaré inequality for upper gradients is proved. Our results indicate that the uniform heat kernel bounds on varying metric measure Dirichlet spaces imply a heat kernel bound on the limit metric measure Dirichlet space. Building on the well-known results on heat kernel estimates [GHL15, Theorem 1.2], Theorem 1.2 in this paper establishes that the package of conditions $\mathrm{(V)}+\mathrm{(PI)}+\mathrm{(CSA)}$ remain stable under Gromov-Hausdorff convergence, where $\mathrm{(V)}$ represents the homogeneous volume growth estimate and $\mathrm{(CSA)}$ represents the cutoff Sobolev inequality (see [AB15] or [GHL15, p.1492] for definition).
@@
This paper is organized as follows. In Section 2, we review the definition of pointed Gromov-Hausdorff convergence. In Section 3, we construct a doubling measure on the limit space and verify Theorem 1.2-(1). In Section 4, we utilize the Hilbert-Schmidt expansion of the heat kernel and estimate the eigenvalues and eigenfunctions. In Section 5, we construct the heat kernel with Dirichlet boundary conditions in the balls of the limit space. In Section 6, we obtain a heat kernel in the limit space and prove Theorem 1.2-(2). In Section 7, the Mosco convergence of Dirichlet forms, stated in Theorem 1.2-(3), is established. In Section 8, we apply Theorem 1.2 to reconstruct the diffusions on the Sierpiński carpet and Sierpiński gasket. We also provide an alternative proof in Theorem 8.4 that for any $\alpha\geq 1$ and $12\leq\beta\leq\alpha+$ 1, there exists a metric measure Dirichlet space that satisfies volume growth $r^{\alpha}$ and $\operatorname{HKE}(r^{\beta})$.
@@ 2 Gromov-Hausdorff convergence
By [Jan17, Lemma 2.8, Proposition 2.9], there are monotone positive sequences $R_{n}\uparrow\infty$ and $s_{n}\downarrow 0$ such that $d_{\text{GH}}(B_{{n}}(p_{{n}},R_{n}),B_{\infty}(p_{\infty},R_{n}))\leq s_{n}\rightarrow 0$ as $n\rightarrow\infty$. For each $n\geq 1$, we apply [HKST15, Proposition 11.1.9] and assume that $(B_{n}(p_{n},R),d_{n})$ and $(B_{\infty}(p_{\infty},R),d_{\infty})$ are isometric embedded into a common metric space $(Z,d_{Z})$. We do not distinguish $(B_{n}(p_{n},R),d_{n})$ and $(B_{\infty}(p_{\infty},R),d_{\infty})$ with their isometric images, respectively. From [HKST15, Proposition 11.1.11 and (11.1.12)] we have
@@
    For each $n\geq 1$ we decompose the volume doubling space $(X_{n},d_{n},m_{n})$ according to Theorem 2.2 and obtain a index set $\{I^{(n)}_{k}\}_{k\in\mathbb{Z}}$, a family of open sets $\left\{Q_{k,j}^{(n)}\subset X_{n}\,\big{|}\,j\in I^{(n)}_{k},k\in\mathbb{Z}\right\}$, a set of points $\left\{z_{k,j}^{(n)}\in X_{n}\,\big{|}\,j\in I^{(n)}_{k},k\in\mathbb{Z}\right\}$ and constants $\delta\in(0,1)$ and $0<a_{0}\leq a_{1}$ that are independent of $n$. Define
@@
    Since $\overline{B_{\infty}(p_{\infty},R)}$ is compact, thus totally bounded, we can find a finite $s_{n}$-net of $\overline{B_{\infty}(p_{\infty},R)}$ with cardinality $N_{n}$, say $\{w_{j}\}_{j=1}^{N_{n}}$. We slightly modify this net so that $\{w_{j}\}_{j=1}^{N_{\epsilon}}\subset{B_{\infty}(p_{\infty},R)}$. For each $z\in\overline{B_{n}(p_{n},R)}$, by (2.9) we are able to find $w_{z}\in\{w_{j}\}_{j=1}^{N_{n}}$ such that $d_{Z}(z,w_{z})\leq 6s_{n}$.
We now define $f_{k,n}:\overline{B_{n}(p_{n},R)}\to\{w_{j}\}_{j=1}^{N_{n}}\cup\{p_{\infty}\}\subset B_{\infty}(p_{\infty},R_{n})$ by
@@
For each $w\in\overline{B_{\infty}(p_{\infty},R_{n})}$, we can find a $v_{w}\in\overline{B_{n}(p_{n},R)}$ such that $d_{Z}(w,v_{w})\leq 6s_{n}$. Since the ball $B_{n}(v_{w},s_{n})$ must intersect one of the set $\{Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\}_{j\in J_{k}^{(n)}}$ as the measure $m_{n}$ has full support, we can choose $\widetilde{v}_{z}\in\bigcup_{j\in J_{k}^{(n)}}\left(Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\right)\setminus\{p_{n}\}$ so that $d_{Z}(w,\widetilde{v}_{w})\leq 7s_{n}$. Therefore
@@ 5 Heat kernel on limit space
    $\left|q^{(n_{k},R)}(t,x,y)-q^{(n_{k},R)}(t,x,y)\right|\lesssim\frac{1}{\delta}\frac{V_{n}(R)}{V_{n_{k}}(\Psi^{-1}_{n_{k}}(\delta/8))^{2}}\absolutevalue{s-t}.$ | . (5.27)
@@ 6 Dirichlet form on limit space
    By the general theory of Dirichlet forms [FOT11, Chapter 1], this Markovian strongly continuous semigroup $\{Q_{t}\}_{t>0}$ corresponds to a unique Dirichlet form, which is denoted by $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$.
@@
    $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ is symmetric and conservative since $\{Q_{t}\}$ is. The heat kernel estimates is given in Theorem 6.5-(5).
@@
    Since $\{Q_{t}\}$ is conservative and admits two-sided heat kernel estimates $\mathrm{HKE}(\Psi_{\infty})$ by Theorem 6.5-(5), we can apply [Lie15, Proposition 3.2] and see that the set $\mathcal{C}:=\{Q_{t}f:f\in C_{0}(X_{\infty})\cap L^{2}(X_{\infty},m_{\infty}),t>0\}$ is a core of $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$. In fact, by [Lie15, Proposition 3.2] and [FOT11, Lemma 1.3.3 (i) and (iii)], we see that $\mathcal{C}\subset C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$, $\mathcal{C}$ is dense in $C_{0}(X_{\infty})$ in the uniform norm and dense in $C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ in $\mathcal{E}_{\infty,1}$. We now prove that $C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ is dense in $\mathcal{F}_{\infty}$ in the $\mathcal{E}_{\infty,1}$ norm, so $\mathcal{C}$ is also dense in $\mathcal{F}_{\infty}$ in the $\mathcal{E}_{\infty,1}$ norm. By [FOT11, Lemma 1.3.3 (iii)], it suffices to prove that $Q_{t}(L^{2}(X_{\infty},m_{\infty}))\subset C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ for each $t>0$.
    From [FOT11, Lemma 1.3.3 (i)], we know that $Q_{t}(L^{2}(X_{\infty},m_{\infty}))\subset\mathcal{F}$. Fix $\epsilon>0$, $f\in L^{2}(X_{\infty},m_{\infty})$ and let $R>0$ large enough such that $\lVert f\rVert_{L^{2}(X_{\infty}\setminus B_{\infty}(p_{\infty},R),m_{\infty})}\leq\epsilon$, then
@@
    By Theorem 6.6-(5) and [Lie15, Proposition 3.2], $\{Q_{t}\}$ or $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ corresponds to a conservative Hunt process on $X_{\infty}$ with continuous path, therefore $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ is strongly local by [FOT11, Theorem 4.5.4].
@@ 7 Mosco convergence of Dirichlet forms
The following definition extends the $L^{2}$ convergence of functions to varying spaces.
@@
Let $\mathcal{E}_{n}$ be a quadratic form on $L^{2}(X_{n},m_{n})$ for each $1\leq n\leq\infty$. Let $\{P_{t}^{(n)}\}_{t>0}$ be their corresponding semigroups. Then $\mathcal{E}_{n}$ Mosco converges to $\mathcal{E}_{\infty}$ if and only if $P_{t}^{(n)}$ converges to $P_{t}^{(\infty)}$ in the strong operator topology, for each $t>0$.
@@
where $\upphi_{n}$ is defined in (3.4). We may assume that the sequences in Proposition 2.3 have the property that $\{\epsilon_{n}\}\subset(0,1/8)$ and $\{R_{j}\}\subset[1,\infty)$. Moreover we may assume that $\mathrm{supp}\left(v_{j}\right)\subset B_{\infty}\left(p_{\infty},R_{j}-1\right)\subset B_{\infty}\left(p_{\infty},R_{j}\right)$ so that $\mathrm{supp}(\upphi_{n}v_{j})\subset B_{n}(p_{n},R_{j})$. Let $\phi_{j}\in C_{c}\left(B_{\infty}\left(p_{\infty},R_{j}\right)\right)$ be a cut-off function such that

### v2

@@ Abstract
We construct a conservative and strongly local regular symmetric Dirichlet form on the pointed Gromov-Hausdorff limit space and demonstrate the stability of heat kernel estimates under this convergence. Furthermore, we establish the Mosco convergence of the associated energy forms along a subsequence.
@@ 1 Introduction
    Since the assumptions and conclusions in Theorem 1.2 are stable under a bi-Lipschitz change of metric, we may instead assume that each $d_{j}$ is (uniformly) bi-Lipschitz equivalent to a length metric, which can be ensured by the quasi-convexity [HKST15, Proposition 8.3.12] or by the chain condition [KM20, Proposition A.1].
@@
We briefly outline the proof of Theorem 1.2. The Radon measure $m_{\infty}$ on limit space is derived from the tightness of the push-forward measures. To obtain the limit Dirichlet form $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$, we first expand the heat kernel of the part of $(\mathcal{E}_{n},\mathcal{F}_{n})$ on the balls $B_{n}(p_{n},R)$ with Dirichlet boundary conditions, using their eigenvalues and eigenfunctions (named Hilbert-Schmidt expansion, see (4.1)). We then apply a version of Arzelà–Ascoli theorem (Lemma 5.1) to take the limit and obtain the limits of eigenvalues and eigenfunctions, which gives a heat kernel on $B_{\infty}(p_{\infty},R)$. Letting $R\to\infty$, we obtain a heat kernel on $X_{\infty}$ (Theorem 6.5). This heat kernel defines a semigroup (Theorem 6.6), corresponding to a Dirichlet form that satisfies the desired properties. The verification of Mosco convergence relies on the strong convergence of semigroups, as stated in the Mosco-Kuwae-Shioya’s theorem (Theorem 7.3).
The construction of limit heat kernels using the Hilbert-Schmidt expansion was first introduced by Ding [Din02] and later studied by Xu [Xu14], in the context of complete Riemannian manifolds with nonnegative Ricci curvature. Their approach relied on gradient estimates and Harnack’s convergence theorem specific to manifolds, which are not applicable in the settings of fractal spaces and general metric measure Dirichlet spaces. In this work, we estimate the eigenvalues and eigenfunctions using only information derived from heat kernel estimates.
@@
In [KS03, Section 5], the authors show that, under the uniform Poincaré inequality and other geometric conditions, the Dirichlet forms converge in the Mosco topology. However, it remains unknown whether the Poincaré inequality $\mathrm{(PI)}$ is stable—i.e., whether it holds for the limit Dirichlet form. In [HKST15, Section 11.6], the stability of the $p$-Poincaré inequality for upper gradients is proved. Our results indicate that the uniform heat kernel bounds on varying metric measure Dirichlet spaces imply a heat kernel bound on the limit metric measure Dirichlet space. Building on the well-known results on heat kernel estimates [GHL15], Theorem 1.2 in this paper establishes that the package of conditions $\mathrm{(V)}+\mathrm{(PI)}+\mathrm{(CSA)}$ should be stable under Gromov-Hausdorff convergence, where $\mathrm{(V)}$ represents the homogeneous volume growth estimate and $\mathrm{(CSA)}$ represents the cutoff Sobolev inequality (see [AB15] or [GHL15, p.1492] for definition).
@@
This paper is organized as follows. In Section 2, we review the definition of pointed Gromov-Hausdorff convergence. In Section 3, we construct a doubling measure on the limit space and verify Theorem 1.2-(1). In Section 4, we utilize the Hilbert-Schmidt expansion of the heat kernel and estimate the eigenvalues and eigenfunctions. In Section 5, we construct the heat kernel with Dirichlet boundary conditions in the balls of the limit space. In Section 6, we obtain a heat kernel in the limit space and prove Theorem 1.2-(2). In Section 7, the Mosco convergence of Dirichlet forms, stated in Theorem 1.2-(3), is established. In Section 8, we apply Theorem 1.2 to reconstruct the diffusions on the Sierpiński carpet and Sierpiński gasket. We also provide an alternative proof in Theorem 8.4 that for any $\alpha\geq 1$ and $12\leq\beta\leq\alpha+$ 1, there exists a metric measure Dirichlet space that satisfies volume growth $r^{\alpha}$ and heat kernel bounds $\operatorname{HKE}_{\mathrm{f}}(r^{\beta})$.
@@ 2 Gromov-Hausdorff convergence
By [Jan17, Lemma 2.8, Proposition 2.9], there are monotone positive sequences $R_{n}\uparrow\infty$ and $s_{n}\downarrow 0$ such that $d_{\text{GH}}(B_{{n}}(p_{{n}},R_{n}),B_{\infty}(p_{\infty},R_{n}))\leq s_{n}\rightarrow 0$ as $n\rightarrow\infty$. For each $n\geq 1$, we apply [HKST15, Proposition 11.1.9] and assume that $(B_{n}(p_{n},R_{n}),d_{n})$ and $(B_{\infty}(p_{\infty},R_{n}),d_{\infty})$ are isometric embedded into a common metric space $(Z,d_{Z})$. We do not distinguish $(B_{n}(p_{n},R),d_{n})$ and $(B_{\infty}(p_{\infty},R_{n}),d_{\infty})$ with their isometric images, respectively. From [HKST15, Proposition 11.1.11 and (11.1.12)] we have
@@
    For each $n\geq 1$, we decompose the volume doubling space $(X_{n},d_{n},m_{n})$ according to Theorem 2.2 and obtain a index set $\{I^{(n)}_{k}\}_{k\in\mathbb{Z}}$, a family of open sets $\left\{Q_{k,j}^{(n)}\subset X_{n}\,\big{|}\,j\in I^{(n)}_{k},k\in\mathbb{Z}\right\}$, a set of points $\left\{z_{k,j}^{(n)}\in X_{n}\,\big{|}\,j\in I^{(n)}_{k},k\in\mathbb{Z}\right\}$, and constants $\delta\in(0,1)$ and $0<a_{0}\leq a_{1}$ that are independent of $n$. Define
@@
    Since $\overline{B_{\infty}(p_{\infty},R_{n})}$ is compact, thus totally bounded, we can find a finite $s_{n}$-net of $\overline{B_{\infty}(p_{\infty},R_{n})}$ with cardinality $N_{n}$, say $\{w_{j}\}_{j=1}^{N_{n}}$. We slightly modify this net so that $\{w_{j}\}_{j=1}^{N_{n}}\subset{B_{\infty}(p_{\infty},R_{n})}$. For each $z\in\overline{B_{n}(p_{n},R_{n})}$, by (2.9) we are able to find $w_{z}\in\{w_{j}\}_{j=1}^{N_{n}}$ such that $d_{Z}(z,w_{z})\leq 6s_{n}$.
We now define $f_{k,n}:\overline{B_{n}(p_{n},R_{n})}\to\{w_{j}\}_{j=1}^{N_{n}}\cup\{p_{\infty}\}\subset B_{\infty}(p_{\infty},R_{n})$ by
@@
For each $w\in\overline{B_{\infty}(p_{\infty},R_{n})}$, we can find a $v_{w}\in\overline{B_{n}(p_{n},R)}$ such that $d_{Z}(w,v_{w})\leq 6s_{n}$. Since the ball $B_{n}(v_{w},s_{n})$ must intersect one of the set $\{Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\}_{j\in J_{k}^{(n)}}$ as the measure $m_{n}$ has full support, we can choose $\widetilde{v}_{w}\in\bigcup_{j\in J_{k}^{(n)}}\left(Q_{k,j}^{(n)}\cap\overline{B_{n}(p_{n},R_{n})}\right)\setminus\{p_{n}\}$ so that $d_{Z}(w,\widetilde{v}_{w})\leq 7s_{n}$. Therefore
@@ 5 Heat kernel on limit space
    $\left|q^{(n_{k},R)}(s,x,y)-q^{(n_{k},R)}(t,x,y)\right|\lesssim\frac{1}{\delta}\frac{V_{n}(R)}{V_{n_{k}}(\Psi^{-1}_{n_{k}}(\delta/8))^{2}}\absolutevalue{s-t}.$ | . (5.27)
@@ 6 Dirichlet form on limit space
To estimate the bounds of $q$, we need the following result on the convergence of $\Phi_{n}$.
@@
    By the general theory of Dirichlet forms [FOT11, Chapter 1], the Markovian strongly continuous semigroup $\{Q_{t}\}_{t>0}$ corresponds to a unique Dirichlet form, which is denoted by $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$.
@@
    $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ is symmetric and conservative since $\{Q_{t}\}$ is. The heat kernel bound is given in Theorem 6.5-(5).
@@
    Since $\{Q_{t}\}$ is conservative and admits two-sided heat kernel estimates $\mathrm{HKE}(\Psi_{\infty})$ by Theorem 6.5-(5), we can apply the proof of [Lie15, Proposition 3.2] and see that the set $\mathcal{C}:=\{Q_{t}f:f\in C_{0}(X_{\infty})\cap L^{2}(X_{\infty},m_{\infty}),t>0\}$ is a core of $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$. In fact, by [Lie15, Proposition 3.2] and [FOT11, Lemma 1.3.3 (i) and (iii)], we see that $\mathcal{C}\subset C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$, $\mathcal{C}$ is dense in $C_{0}(X_{\infty})$ in the uniform norm and dense in $C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ in $\mathcal{E}_{\infty,1}$. We now prove that $C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ is dense in $\mathcal{F}_{\infty}$ in the $\mathcal{E}_{\infty,1}$ norm, so $\mathcal{C}$ is also dense in $\mathcal{F}_{\infty}$ in the $\mathcal{E}_{\infty,1}$ norm. By [FOT11, Lemma 1.3.3 (iii)], it suffices to prove that $Q_{t}(L^{2}(X_{\infty},m_{\infty}))\subset C_{0}(X_{\infty})\cap\mathcal{F}_{\infty}$ for each $t>0$.
    From [FOT11, Lemma 1.3.3 (i)], we know that $Q_{t}(L^{2}(X_{\infty},m_{\infty}))\subset\mathcal{F}_{\infty}$. Fix $\epsilon>0$, $f\in L^{2}(X_{\infty},m_{\infty})$ and let $R>0$ large enough such that $\lVert f\rVert_{L^{2}(X_{\infty}\setminus B_{\infty}(p_{\infty},R),m_{\infty})}\leq\epsilon$, then
@@
    Let $f,g\in\mathcal{F}_{\infty}$, $K$ be a compact set such that $\mathrm{supp}(g)\subset K$ and assume that $f$ is constant on an open set $U$ that contains $K$. By [FOT11, (4.5.7)] and the conservativeness of $\{Q_{t}\}$, we have
    $\mathcal{E}_{\infty}(f,g)$ $=\lim_{t\to 0}\frac{1}{2t}\int_{X_{\infty}}\int_{X_{\infty}}(f(x)-f(y))(g(x)-g(y))q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(x)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.34)
    $=\lim_{t\to 0}\frac{1}{2t}\Big{(}\int_{X_{\infty}\setminus U}\int_{K}(f(x)-f(y))(g(x)-g(y))q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(x)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.35)
    $\quad+\int_{K}\int_{X_{\infty}\setminus U}(f(x)-f(y))(g(x)-g(y))q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(x)\mathop{}\!\mathrm{d}m_{\infty}(y)\Big{)}.$ (6.36)
    By the bound of $q$, we have for $\delta:=\mathrm{dist}(K,X_{\infty}\setminus U)$ that
    $\phantom{\leq\ }\int_{X_{\infty}\setminus U}\int_{K}(f(x)-f(y))(g(x)-g(y))q(t,x,y)\mathop{}\!\mathrm{d}m_{\infty}(x)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.37)
    $\lesssim\frac{1}{V_{\infty}(\Psi_{\infty}^{-1}(t))}\exp(-c\left(\frac{\Psi_{\infty}(\delta)}{t}\right)^{1/(\beta^{\prime}-1)})\int_{X_{\infty}\setminus U}\int_{K}(f(x)-f(y))(g(x)-g(y))\mathop{}\!\mathrm{d}m_{\infty}(x)\mathop{}\!\mathrm{d}m_{\infty}(y)$ (6.38)
    $\to 0\quad(t\to 0)$ (6.39)
    Similarly we can estimate the second term in (6.36). Therefore $\mathcal{E}_{\infty}(f,g)=0$ and $(\mathcal{E}_{\infty},\mathcal{F}_{\infty})$ is strongly local.
@@ 7 Mosco convergence of Dirichlet forms
The following definition, from [KS03], extends the $L^{2}$ convergence of functions to varying spaces.
@@
Let $\mathcal{E}_{n}$ be a quadratic form on $L^{2}(X_{n},m_{n})$ for each $1\leq n\leq\infty$. Let $\{P_{t}^{(n)}\}_{t>0}$, $1\leq n\leq\infty$ be their corresponding semigroups. Then $\mathcal{E}_{n}$ Mosco converges to $\mathcal{E}_{\infty}$ if and only if $P_{t}^{(n)}$ converges to $P_{t}^{(\infty)}$ in the strong operator topology, for each $t>0$.
@@
where $\upphi_{n}$ is defined in (3.4). We may assume that the sequences in Proposition 2.3 satisfying $\{\epsilon_{n}\}\subset(0,1/8)$ and $\{R_{j}\}\subset[1,\infty)$. Moreover we may assume that $\mathrm{supp}\left(v_{j}\right)\subset B_{\infty}\left(p_{\infty},R_{j}-1\right)\subset B_{\infty}\left(p_{\infty},R_{j}\right)$ so that $\mathrm{supp}(\upphi_{n}v_{j})\subset B_{n}(p_{n},R_{j})$. Let $\phi_{j}\in C_{c}\left(B_{\infty}\left(p_{\infty},R_{j}\right)\right)$ be a cut-off function such that